OpenStudy (anonymous):

I really need help on this question. I have exhausted all of my options... I do not just want an answer I would like it explained to me. Which of the following statements have the same result? Explain each step in solving each one. f(1) when f(x) = 5x + 1 f-1(3) when f(x) = "2x plus 3, all over 5 3y - 7 = y + 5

OpenStudy (mathmale):

f-1(3) when f(x) = "2x plus 3, all over 5 is a bit ambiguous. Did you mean the inverse function, \[f ^{-1}(3) = \frac{ 2x+3 }{ 5 }?\]

OpenStudy (mathmale):

Ensure that nothing is missing from your typing in the problem statement and possible answers.

OpenStudy (anonymous):

Yes that is what I meant. Sorry.

OpenStudy (anonymous):

Everything else is correct. I apologize.

OpenStudy (mathmale):

f(1) when f(x) = 5x + 1 is relatively easy. Just substitute 1 for x on the right side of this equation. 3y - 7 = y + 5 can be solved for y. What would be the resulting value of y? The remaining relationship is problematic because on one hand, you're stating that x=5 on the left side, and on the other hand x has not been replaced with 5. I'd double-check that you've typed this relationship correctly. Hint: Try solving for f(1) in the first line. Then try solving for y in the second line. What do you see?

OpenStudy (anonymous):

Okay so for the first one I got 6....

OpenStudy (anonymous):

and for the other one I typed it correctly... but I am sort of confused on how I set this one up to solve it: 3y-7=y+5...

OpenStudy (anonymous):

Oh wait, it is just simple algebra... that one is also 6..

OpenStudy (anonymous):

@mathmale ... can you please explain to me how to do \[f^-1(3)=(2x+3)/(5)\]

OpenStudy (mathmale):

The last problem really is not a solvable problem. Note how you obtained 6 for each of the first two. What does that tell you about the answer for this question?

OpenStudy (anonymous):

That those two have the same result. So that would be the answer. And then I just explain how to do them.

OpenStudy (anonymous):

Thank you so much! (: I guess I knew how to do it, I just needed to be pointed in the right direction.

OpenStudy (mathmale):

:)